# Cross-correlation scheme for quantum optical coherence tomography based on Michelson interferometer

### v1

@@ III Experiment
To support our calculations, we conducted the following experiment (Fig. 4). A degenerate SPDC source (see details in Appendix B) with a $50$ nm bandwidth and a central wavelength of $1313$ nm was coupled into a Michelson interferometer built with a fiberized beam-splitter (FBS). The reference and sample arms were taken out into free space. The mirror in the sample arm was translated with piezo- and stepper-motor stage to record interferograms.
To record both cross-correlation $M_{\text{MI}}^{\text{(cross)}}(\tau)$ and autocorrelation $M_{\text{MI}}^{\text{(auto)}}(\tau)$ interferograms simultaneously, we inserted a fiber circulator at the interferometer input and added a second FBS at the output. Photon detection was performed with three near-infrared single‑photon detectors (ID Qube NIR), and coincidences were registered using a Time Tagger Ultra (Swabian Instruments). Autocorrelation interferograms were obtained from D2–D3 coincidences, while cross‑correlation interferograms were derived by summing D1–D2 and D1–D3 coincidences.
@@
  • Defectoscopy application: Two silicon windows served as a dispersive sample, and we resolved the air gap between them. To compensate for potential non‑parallelism of the window surfaces the SPDC radiation was focused by the lens L. Also it allows transverse scan of the sample as in 3D-imaging QOCT system.
@@ IV Experimental results and discussion
Figure 5: Experimental results: (a) Single-count PSF interferogram for OCT: raw data (purple), envelope (thick bright-red dashed), theoretical envelope (thin dark-red dashed). (b) Coincidence-count PSF interferograms for QOCT: raw cross-correlation ($M_{\mathrm{MI}}^{\mathrm{(cross)}}$, pale red) and autocorrelation ($M_{\mathrm{MI}}^{\mathrm{(auto)}}$, pale blue), filtered HOM dips (solid dark red and dark blue), Gaussian fits (red and blue dashed). (c) Absolute FFT of the OCT interferogram in (a): raw (purple), envelope (thick bright-red dashed), SPDC source spectrum (thin dark-red dashed). (d) FFTs of the QOCT data: solid red and blue curves, with darker segments indicating the filter region; dashed lines are Gaussian envelopes. (e,f) Single-count OCT and coincidence-count QOCT PSF interferograms with dispersion introduced by a 5 mm silicon window. (g,h) Single-count OCT and coincidence-count QOCT interferograms for a sample of two silicon windows separated by a $\sim$100 $\upmu\mathrm{m}$ air gap.
@@ IV.1 Proof-of-concept experiment
The results of the proof-of-concept experiment are shown in Fig. 5(a–d). To compare QOCT with standard OCT resolution we measured a single-count interferogram at one interferometer output, shown as a purple line in Fig. 5a. A sinc fit to its envelope coincides with the theoretical envelope – computed as the FFT of the measured source spectrum (dark-red dashed curve in Fig. 5a) – with FWHM values of $16.6\pm 0.1\ \upmu$m and $16.4\pm 0.1\ \upmu$m, respectively. Here and below, uncertainties are fitting errors. Despite this agreement, the sinusoidal carrier exhibits significant phase noise from fiber fluctuations, which manifests as broadening in the Fourier domain.
Figure 5c shows the FFT magnitude of the OCT interferogram. A Gaussian fit to its envelope yields $\mathrm{FWHM}=0.39\pm 0.02\,\omega_{\mathrm{p}}$, whereas the source spectrum itself (dark-red dashed) has $\mathrm{FWHM}=0.02\,\omega_{\mathrm{p}}$. Thus, phase instability broadens the OCT interferogram so much that Fourier filtering alone would be challenging.
The raw QOCT coincidence interferograms $M_{\mathrm{MI}}^{\mathrm{(cross)}}(\tau)$ (pale red) and $M_{\mathrm{MI}}^{\mathrm{(auto)}}(\tau)$ (pale blue) are plotted in Fig. 5b. These were recorded with a 70 nm step size and 100 ms exposition per point.
Their FFT magnitudes appear in Fig. 5d. The peak at $\omega_{\mathrm{p}}/2$ (the $\tilde{M}_{1}$ term) is present in the autocorrelation spectrum but vanishes in the cross-correlation spectrum; its Gaussian envelope has $\mathrm{FWHM}=0.61\pm 0.05\,\omega_{\mathrm{p}}$, causing some overlap with the desired $\tilde{M}_{0}$ peak. Conversely, the peak at $\omega_{\mathrm{p}}$ (the $\tilde{M}_{2}$ term) is clear in the cross-correlation spectrum but obscured in the autocorrelation one, broadened by phase fluctuations to $\mathrm{FWHM}=0.94\pm 0.05\,\omega_{\mathrm{p}}$.
To extract the HOM term $\tilde{M}_{0}$, we discarded FFT components above a cutoff of $0.015\,\omega_{\mathrm{p}}$ and applied an inverse FFT. The processed interferograms (thick red and blue curves in Fig. 5b) have FWHMs of $9.39\pm 0.25\ \upmu$m and $8.42\pm 0.35\ \upmu$m, respectively – roughly half the OCT PSF. The slight discrepancy may arise because the single-count spectrum is broader than the coincidence-count spectrum due to source coupling into the single-mode fiber [7].
@@ IV.2 Dispersion cancellation test
The results of the dispersion cancellation experiment are shown in Fig. 5(e,f). To demonstrate the effect of the 5 mm silicon window on OCT, we first measured the single-count interferogram (OCT PSF) in Fig. 5e. Its envelope broadened more than tenfold, to a FWHM of $110.9\pm 6.8\leavevmode\nobreak\ \upmu$m.
The raw QOCT interferograms, recorded with a 70 nm step size and 200 ms exposition per point, are shown in Fig. 5f. We extracted the HOM terms as in the previous section. The FWHM of the processed QOCT PSFs remains nearly unchanged: $9.52\pm 0.07\leavevmode\nobreak\ \upmu$m for cross‐correlation and $14.31\pm 0.39\leavevmode\nobreak\ \upmu$m for autocorrelation. However, the silicon window introduces additional insertion losses, reducing the coincidence count rates to $0.175$ kcps (cross‐correlation) and $0.051$ kcps (autocorrelation), which made fitting the autocorrelation data challenging.
These results confirm dispersion cancellation in both schemes; however, the cross‐correlation scheme is more robust to loss thanks to its higher signal-to-noise ratio and simpler spectral filtering. Therefore, we employ only the cross‐correlation scheme in the following defectoscopy application.
@@ IV.3 Defectoscopy application
In the previous experiments, we compared the two QOCT measurement schemes using sub-wavelength scans on a piezo translator to enable post-processing via Fourier filtering. However, this approach does not utilize the key advantage of HOM interference – robustness to optical phase fluctuations and no requirement for sub-wavelength path-length control – and is limited by the piezo’s small travel range.
For the defectoscopy application, we therefore recorded interferograms by continuously moving the mirror with a stepper-motor translator at a constant speed of 16.7 nm/s and summing counts every 300 nm to average over high-frequency phase noise. Each scan was repeated five times to improve precision, and the stepper-motor’s large range enabled us to probe the thick sample.
Figures 5(g,h) show the measured interferograms for an air gap between two silicon windows. The single-photon OCT interferogram in Fig. 5g is severely blurred by dispersion, making layer separation impossible. In contrast, the QOCT coincidence interferogram in Fig. 5h exhibits distinct peaks and dips – corresponding to reflections, echoes, and cross-interference terms [17] – allowing us to resolve the window separation. Appendix C provides a detailed description of each enumerated term and the reconstruction algorithm. From this analysis, we extract a gap width of $110.67\pm 0.11\leavevmode\nobreak\ \upmu\mathrm{m}$, in excellent agreement with the known spacer thickness.
@@ V Conclusion
We have studied a cross‑correlation QOCT scheme based on a Michelson interferometer that overcomes the main limitations of previous implementations while preserving all of QOCT’s advantages. It maintains an axial resolution twice that of classical OCT for a given source spectrum and demonstrates full dispersion cancellation, even when OCT resolution is broadened by more than tenfold.
Unlike the standard autocorrelation scheme, it achieves a four-fold higher signal-to-noise ratio and the interferogram’s spectrum contains only a single parasitic peak at the pump frequency, far from the target HOM‑peak at zero. Although this extra peak is ideally narrowband, in practice it can broaden significantly due to phase instability; nevertheless, its large spectral separation from the HOM term enables straightforward spectral filtering without limiting source bandwidth, allowing a fiberized implementation that is compact and vibration-insensitive. Furthermore, the desired HOM term can be extracted simply by averaging over high‑frequency oscillations during continuous scanning with a coarse mechanical translator with a large range, greatly simplifying the scan process. These features make the technique suitable for real-world applications such as defectoscopy, where it enables precise crack localization in highly dispersive media.
@@ Acknowledgements.
Appendix A Photon pair interference in a Michelson interferometer: general description
Denote the sample response function as $H(\omega)$ and delay in the reference channel as $\tau$. Then input and output modes are connected as follows:
Combining equations (1) and (2), we obtain:
$\hat{a}_{\text{in}}^{\dagger}=\frac{1}{2}\left[H\left(\omega\right)+{{\operatorname{e}}^{\operatorname{i}\omega\tau}}\right]\hat{a}_{\text{out}}^{\dagger}+\frac{\operatorname{i}}{2}\left[H\left(\omega\right)-{{\operatorname{e}}^{\operatorname{i}\omega\tau}}\right]\hat{b}_{\text{out}}^{\dagger}.$ (12)
Assume $f\left({{\omega}_{s}},{{\omega}_{i}}\right)$ is symmetrical, so that $f\left({{\omega}_{s}},{{\omega}_{i}}\right)=f\left({{\omega}_{i}},{{\omega}_{s}}\right)$, which is true for type-I or type-0 SPDC.
@@ Appendix A Photon pair interference in a Michelson interferometer: general description
$\left|{{\Psi}_{\text{in}}}\right\rangle=\frac{1}{{4\sqrt{2}}}\times$ (13)
@@
$\operatorname{i}$ $\iint{\text{d}{{\omega}_{s}}\text{d}{{\omega}_{i}}}f_{\text{out}}^{(ab)}\left({{\omega}_{s}},{{\omega}_{i}}\right)\hat{a}^{\dagger}_{\text{out}}(\omega_{\text{s}})\hat{b}^{\dagger}_{\text{out}}(\omega_{\text{i}})\left|\text{vac}\right\rangle-$
$\iint{\text{d}{{\omega}_{s}}\text{d}{{\omega}_{i}}}f_{\text{out}}^{(bb)}\left({{\omega}_{s}},{{\omega}_{i}}\right)\hat{b}^{\dagger}_{\text{out}}(\omega_{\text{s}})\hat{b}^{\dagger}_{\text{out}}(\omega_{\text{i}})\left|\text{vac}\right\rangle\Bigg{]}$
@@
$f_{\text{out}}^{(aa)}\left({{\omega}_{s}},{{\omega}_{i}}\right)\equiv$ $f\left({{\omega}_{s}},{{\omega}_{i}}\right)\times$
$\left[H(\omega_{\text{s}})+\operatorname{e}^{\operatorname{i}\omega_{\text{s}}\tau}\right]\left[H(\omega_{\text{i}})+\operatorname{e}^{\operatorname{i}\omega_{\text{i}}\tau}\right],$ (14a)
$f_{\text{out}}^{(ab)}\left({{\omega}_{s}},{{\omega}_{i}}\right)\equiv$ $2f\left({{\omega}_{s}},{{\omega}_{i}}\right)\times$
$\left[H(\omega_{\text{s}})+\operatorname{e}^{\operatorname{i}\omega_{\text{s}}\tau}\right]\left[H(\omega_{\text{i}})-\operatorname{e}^{\operatorname{i}\omega_{\text{i}}\tau}\right],$ (14b)
$f_{\text{out}}^{(bb)}\left({{\omega}_{s}},{{\omega}_{i}}\right)\equiv$ $f\left({{\omega}_{s}},{{\omega}_{i}}\right)\times$
$\left[H(\omega_{\text{s}})-\operatorname{e}^{\operatorname{i}\omega_{\text{s}}\tau}\right]\left[H(\omega_{\text{i}})-\operatorname{e}^{\operatorname{i}\omega_{\text{i}}\tau}\right].$ (14c)
The projections are given by
$\quad\frac{1}{8}\left[{{f}^{(bb)}_{\text{out}}}\left({{\omega}_{1}},{{\omega}_{2}}\right)+{{f}^{(bb)}_{\text{out}}}\left({{\omega}_{2}},{{\omega}_{1}}\right)\right]=$
$\quad\frac{1}{4}f\left({{\omega}_{1}},{{\omega}_{2}}\right)\left[H(\omega_{1})-\operatorname{e}^{\operatorname{i}\omega_{1}\tau}\right]\left[H(\omega_{2})-\operatorname{e}^{\operatorname{i}\omega_{2}\tau}\right],$
$\left\langle{{\Psi}^{(ab)}_{M}}\left({{\omega}_{1}},{{\omega}_{2}}\right)|{{\Psi}_{\text{out}}}\right\rangle=\frac{1}{2\sqrt{2}}{{f}^{(ab)}_{\text{out}}}\left({{\omega}_{1}},{{\omega}_{2}}\right)=$ (15b)
$\quad\frac{1}{2\sqrt{2}}f\left({{\omega}_{1}},{{\omega}_{2}}\right)\left[H(\omega_{1})+\operatorname{e}^{\operatorname{i}\omega_{1}\tau}\right]\left[H(\omega_{2})-\operatorname{e}^{\operatorname{i}\omega_{2}\tau}\right].$
Let’s make a change of variables: $\omega_{1}=\omega_{0}+\nu_{1}$, $\omega_{2}=\omega_{0}+\nu_{2}$, where $\omega_{0}\equiv\omega_{\text{p}}/2$ – half of the pump frequency. In this case, we have
$\Big{|}\left[H(\omega_{0}+\nu_{1})-\operatorname{e}^{\operatorname{i}(\omega_{0}+\nu_{1})\tau}\right]$$\times$ (16a)
$\left[H(\omega_{0}+\nu_{2})-\operatorname{e}^{\operatorname{i}(\omega_{0}+\nu_{2})\tau}\right]$ $\Big{|}^{2}=$
$K_{c}+K_{0}$ $-(K_{1}+K_{1^{\prime}})/2+K_{2},$
$\Big{|}\left[H(\omega_{0}+\nu_{1})+\operatorname{e}^{\operatorname{i}(\omega_{0}+\nu_{1})\tau}\right]$$\times$ (16b)
$\left[H(\omega_{0}+\nu_{2})-\operatorname{e}^{\operatorname{i}(\omega_{0}+\nu_{2})\tau}\right]$ $\Big{|}^{2}=$
$K_{c}-K_{0}$ $+(K_{1}-K_{1^{\prime}})/2-K_{2},$
where
$K_{c}\equiv$ (|$\left(\absolutevalue{H(\omega_{0}+\nu_{1})}^{2}+1\right)\left(\absolutevalue{H(\omega_{0}+\nu_{2})}^{2}+1\right),$ (17a)
@@
$16M^{(bb)}(\tau)=M_{c}^{(M)}+M_{0}(\tau)-\frac{M_{1}(\tau)+M_{1^{\prime}}(\tau)}{2}+M_{2}(\tau),$ (18a)
$8M^{(ab)}(\tau)=M_{c}^{(M)}-M_{0}(\tau)+\frac{M_{1}(\tau)-M_{1^{\prime}}(\tau)}{2}-M_{2}(\tau),$ (18b)
@@
for $j=c,0,1,1^{\prime},2$. It is easy to see that $M_{1}(\tau)=M_{1^{\prime}}(\tau)$, therefore finally we have the Eqs. (9).

### v2

@@ III Experiment
To support our calculations, we conducted an experiment to benchmark the studied cross-correlation MI-QOCT against both the autocorrelation scheme and classical OCT, and to assess its potential applicability to real-world scenarios. The experimental setup is presented in Fig. 4. A degenerate SPDC source (see details in Appendix B) with a $50$ nm bandwidth and a central wavelength of $1313$ nm was coupled into a Michelson interferometer built with a fiberized beam-splitter (FBS). The reference and sample arms were taken out into free space. The mirror in the sample arm was translated with piezo or a stepper-motor stage to record interferograms.
To record both cross-correlation $M_{\text{MI}}^{\text{(cross)}}(\tau)$ and autocorrelation $M_{\text{MI}}^{\text{(auto)}}(\tau)$ interferograms simultaneously, we inserted a fiber circulator at the interferometer input and added a second FBS at the output. Photon detection was performed with three near-infrared single‑photon detectors (ID Qube NIR), and coincidences were registered using a Time Tagger Ultra (Swabian Instruments). Autocorrelation interferograms were obtained from D2–D3 coincidences, while cross‑correlation interferograms were derived by summing D1–D2 and D1–D3 coincidences. Additionally, single-photon interferograms, related to classical OCT, were obtained from the sum of D2 and D3 count rates.
@@
  • Defectoscopy application: Two silicon windows served as a dispersive sample, and we resolved the air gap between them. To compensate for potential non‑parallelism of the window surfaces, the SPDC radiation was focused by lens L. This also potentially allows transverse scanning of the sample, as in a 3D-imaging QOCT system.
@@ IV Experimental results and discussion
Figure 5: Experimental results: (a) Single-count PSF interferogram for OCT: raw data (light-purple), envelope fit (thick dark-purple dashed), theoretical envelope (thick brown) obtained by inverse FFT of the SPDC source spectrum. (b) Coincidence-count PSF interferograms for QOCT: raw cross-correlation ($M_{\mathrm{MI}}^{\mathrm{(cross)}}$, thin light-blue) and autocorrelation ($M_{\mathrm{MI}}^{\mathrm{(auto)}}$, thin light-red), filtered HOM dips (thick red and blue), Gaussian fits (thick red and blue dashed). (c) Absolute FFT of the OCT interferogram in (a): raw (light-purple), envelope fit (thick dark-purple dashed), SPDC source spectrum (thick brown). (d) FFTs of the QOCT data: solid blue and red curves, with darker segments indicating the filter region; dashed lines are Gaussian envelopes. (e,f) Single-count OCT and coincidence-count QOCT PSF interferograms with dispersion introduced by a 5 mm silicon window. (g,h) Single-count OCT and coincidence-count QOCT interferograms for a sample of two silicon windows separated by a $\sim$110 $\upmu\mathrm{m}$ air gap.
@@ IV.1 Proof-of-concept experiment
The results of the proof-of-concept experiment are shown in the light-green block of Fig. 5.
[…]
The FFT magnitude of this interferogram is shown in Fig. 5c. A Gaussian fit to its envelope (thick dark-purple dashed curve) yielded a spectral FWHM of $0.39\pm 0.02\,\omega_{\mathrm{p}}$.
For comparison, the directly measured SPDC spectrum (already shown in the bottom of Fig. 4) is plotted as a brown line in Fig. 5c, and its inverse FFT is shown as the same line in Fig. 5a. We see that the interferogram envelope matches perfectly with the inverse FFT of the SPDC spectrum (as expected theoretically). However, the interferogram spectrum is nearly 20 times broader than the SPDC spectrum, indicating strong broadening due to phase noise in the fibers. Thus, although the SPDC spectrum fully determines the OCT resolution, the interferogram bandwidth is dominated by phase fluctuations.
The raw QOCT coincidence interferograms $M_{\mathrm{MI}}^{\mathrm{(cross)}}(\tau)$ (light blue) and $M_{\mathrm{MI}}^{\mathrm{(auto)}}(\tau)$ (light red) are shown in Fig. 5b. These were recorded with a 70 nm step size and 100 ms exposure per point.
Their FFT magnitudes are plotted in Fig. 5d with the same color coding. The envelopes were fitted with sums of Gaussian functions (dashed lines) centered at $\omega_{\mathrm{p}}/2$ and $\omega_{\mathrm{p}}$ (see details in Appendix D.1). As expected, the autocorrelation spectrum $\tilde{M}_{\mathrm{MI}}^{\mathrm{(auto)}}$ exhibits both parasitic peaks $\tilde{M}_{1}$ and $\tilde{M}_{2}$. However, in contrast to the theory (Fig. 1c), the $\tilde{M}_{2}$ peak is not narrow but instead twice as broad as $\tilde{M}_{1}$. This can be explained as follows: the same amplitude of path fluctuation produces different phase modulations at different wavelengths, resulting in a spectral broadening of $\tilde{M}_{2}$ that is roughly twice that of $\tilde{M}_{1}$.
These results, shown in Fig. 5d, demonstrate that even when the SPDC spectrum is relatively narrow and the parasitic peaks are not expected to overlap with the HOM term $\tilde{M}_{0}$, phase instability can still broaden them significantly, complicating digital spectral filtering.
By contrast, as expected, the $\tilde{M}_{\mathrm{MI}}^{\mathrm{(cross)}}(\tau)$ spectrum shows only the $\tilde{M}_{2}$ parasitic peak (the fitted amplitude of $\tilde{M}_{1}$ is negligible).
[…]
The small discrepancy likely arises because the single-count spectrum is broader than the coincidence-count spectrum due to coupling into the single-mode fiber [okanoDispersionCancellationHighresolution2013].
@@ IV.2 Dispersion cancellation test
[…]
These results confirm dispersion cancellation in both schemes; however, the cross‐correlation scheme is more robust to loss due to its higher SNR and simpler spectral filtering.
To verify that the cross-correlation scheme’s SNR is suitable for real-world applications, we have conducted a demonstrative defectoscopy experiment, described below.
@@ IV.3 Defectoscopy application
In the previous experiments, we compared the two QOCT measurement schemes by performing sub‑wavelength, step‑scan measurements on a piezotranslator. This fine scanning allowed us to capture the full interferogram, including high‑frequency fringes, and then apply a digital low‑pass filter in post‑processing to isolate the low‑frequency HOM term. However, this method neither exploits the intrinsic phase‑noise insensitivity of HOM interference – since it still requires sub‑wavelength path‑length control – nor overcomes the piezo’s limited travel (100 µm).
For the defectoscopy application, we switched to hardware filtering. We drove a stepper‑motor translator continuously at 16.7 nm/s and binned photon counts into 300 nm intervals, which averages out the high‑frequency fringes in real time. Repeating each scan five times improved precision, and the stepper motor’s 25 mm range allowed us to interrogate the full thickness of the dispersive sample without sub‑wavelength positioning.
Figs 5(g,h) show the measured interferograms for an air gap between two silicon windows.
Both OCT and QOCT interferograms were measured in parallel. To improve the SNR, the scan was repeated 8 times.
The single-photon OCT interferogram in Fig. 5g is severely blurred by dispersion, making layer separation impossible. In contrast, the QOCT coincidence interferogram in Fig. 5h exhibits distinct peaks and dips – corresponding to reflections, echoes, and cross-interference terms [li-gomezQuantumEnhancedProbing2023] – allowing us to resolve the window separation. Appendix D.3 provides a detailed description of each enumerated term and the reconstruction algorithm. From this analysis, we extract a gap width of $110.67\pm 0.11\penalty 10000\ \upmu\mathrm{m}$, in excellent agreement with the known spacer thickness.
@@ V Conclusion
We have studied a cross‑correlation QOCT scheme based on a Michelson interferometer that overcomes the main limitations of previous implementations while preserving all of QOCT’s advantages. Compared with classical OCT, our QOCT scheme delivers a two-fold narrowing of the PSF and remains essentially immune to the sample dispersion that increases the classical OCT PSF by more than a factor of six.
Compared to the standard autocorrelation QOCT, the cross-correlation measurements achieve a fourfold improvement in SNR. Achieving high axial resolution requires a broadband light source. However, the autocorrelation scheme may not fully benefit from this, as it contains a parasitic peak at half the pump frequency, whose width increases with the source bandwidth. This peak overlaps with the HOM term, making it difficult to extract useful information. A cross-correlation QOCT’s interferogram spectrum contains only a single parasitic peak at the pump frequency – well separated from the HOM peak – which greatly simplifies extraction of the target HOM term by post-processing. This spectral separation reduces sensitivity to phase-instability-induced broadening and allows a compact, fiberized, vibration-insensitive implementation. Moreover, this spectral separation potentially enables combining our cross-correlation scheme with ultra-broadband SPDC sources to approach sub-wavelength axial resolution [katamadze_broadband_2025].
We have also shown that the HOM term can be directly recovered by averaging over the high-frequency oscillations during a continuous coarse mechanical scan, obviating the need for sub-wavelength path-length control and enabling long-range scanning. Using this approach, we have shown that the SNR of our technique is high enough to localize a $110\penalty 10000\ \upmu\mathrm{m}$ air gap in a highly dispersive silicon medium.
Thus, the presented QOCT scheme enables the use of bright, broadband collinear type-0/type-I SPDC sources, is immune to optical dispersion and phase instability, and produces clear HOM dips with improved SNR – paving the way to real-world QOCT applications.
@@ Acknowledgements.
Appendix A Single-photon and photon pair interference in a Michelson interferometer: general description
Denote the sample response function as $H(\omega)$ and delay in the reference channel as $\tau$. The input and output modes of the interferometer are related according to Eqs. (1, 2). Combining these equations, we obtain:
$\hat{a}_{\text{in}}^{\dagger}(\omega)=\alpha(\omega)\hat{a}_{\text{out}}^{\dagger}(\omega)+\beta(\omega)\hat{b}_{\text{out}}^{\dagger}(\omega).$ (14a)
were
$\alpha(\omega)\equiv\frac{{{\operatorname{e}}^{\operatorname{i}\omega\tau}}+H\left(\omega\right)}{2},\quad\beta(\omega)\equiv\operatorname{i}\frac{{{\operatorname{e}}^{\operatorname{i}\omega\tau}}-H\left(\omega\right)}{2}.$ (14b)
A.1 Single-photon interference
Consider a general case of a mixed single-photon state in the input mode $a$:
[…]
were $\rho(\omega,\omega^{\prime})$ is a cross-spectral density function.
Single-photon measurement in the mode $b$ corresponds to a projection to a state:
[…]
Then the total probability to detect photon in the output mode $b$ depending on the delay $\tau$ gives the interferogram:
[…]
Substituting Eqs. (1416) into Eq. (17) we obtain:
[…]
Note that Eq. (18) depends on the diagonal part of the crossed-spectral density function only, which describes the spectral distribution of the state and does not depend on its coherence and purity. So, the same interferogram could be obtained with coherent or thermal light used in classical OCT setups.
For the Gaussian spectral distribution (4) and the single-layer response function (3) the Eq. (18) transforms to Eq. (5).
A.2 Photon-pair interference
Consider a photon-pair input state (6) in the input mode $a$. Assume the spectral amplitude $F\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)$ is symmetrical, so that $F\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)=F\left({{\omega}_{\text{i}}},{{\omega}_{\text{s}}}\right)$, which is true for type-I or type-0 SPDC.
@@ A.2 Photon-pair interference
$\left|{{\Psi}_{\text{in}}}\right\rangle=\frac{1}{{\sqrt{2}}}\times$ (19)
$\Bigg[$ $\iint{\text{d}{{\omega}_{\text{s}}}\text{d}{{\omega}_{\text{i}}}}F_{\text{out}}^{(aa)}\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\hat{a}^{\dagger}_{\text{out}}(\omega_{\text{s}})\hat{a}^{\dagger}_{\text{out}}(\omega_{\text{i}})\left|\text{vac}\right\rangle+$
$\iint{\text{d}{{\omega}_{\text{s}}}\text{d}{{\omega}_{\text{i}}}}F_{\text{out}}^{(ab)}\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\hat{a}^{\dagger}_{\text{out}}(\omega_{\text{s}})\hat{b}^{\dagger}_{\text{out}}(\omega_{\text{i}})\left|\text{vac}\right\rangle+$
$\iint{\text{d}{{\omega}_{\text{s}}}\text{d}{{\omega}_{\text{i}}}}F_{\text{out}}^{(bb)}\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\hat{b}^{\dagger}_{\text{out}}(\omega_{\text{s}})\hat{b}^{\dagger}_{\text{out}}(\omega_{\text{i}})\left|\text{vac}\right\rangle\Bigg]$
were
$F_{\text{out}}^{(aa)}\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\equiv$ $\alpha(\omega_{\text{s}})\alpha(\omega_{\text{i}})F\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)$ (20a)
$F_{\text{out}}^{(ab)}\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\equiv$ $2\alpha(\omega_{\text{s}})\beta(\omega_{\text{i}})F\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)$ (20b)
$F_{\text{out}}^{(bb)}\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)\equiv$ $\beta(\omega_{\text{s}})\beta(\omega_{\text{i}})F\left({{\omega}_{\text{s}}},{{\omega}_{\text{i}}}\right)$
The projections for autocorrelation and cross-correlation measurements are given by
$\quad\frac{1}{2}\left[{F^{(bb)}_{\text{out}}}\left({{\omega}_{1}},{{\omega}_{2}}\right)+{F^{(bb)}_{\text{out}}}\left({{\omega}_{2}},{{\omega}_{1}}\right)\right]=$
$\quad\beta(\omega_{1})\beta(\omega_{2})F(\omega_{1},\omega_{2}),$
$\left\langle{{\Psi}^{(ab)}_{M}}\left({{\omega}_{1}},{{\omega}_{2}}\right)|{{\Psi}_{\text{out}}}\right\rangle=\frac{1}{\sqrt{2}}{F}^{(ab)}_{\text{out}}\left({{\omega}_{1}},{{\omega}_{2}}\right)=$ (21b)
$\quad\frac{1}{\sqrt{2}}\alpha(\omega_{1})\beta(\omega_{2})F(\omega_{1},\omega_{2})$
respectively.
Let us make a change of variables: $\omega_{1}=\omega_{0}+\nu_{1}$, $\omega_{2}=\omega_{0}+\nu_{2}$, where $\omega_{0}\equiv\omega_{\text{p}}/2$ – half of the pump frequency. In this case, we have
@@
$\quad\quad K_{c}+K_{0}-(K_{1}+K_{1^{\prime}})/2+K_{2},$
[…]
$\quad\quad K_{c}-K_{0}+(K_{1}-K_{1^{\prime}})/2-K_{2},$
@@
[…]
@@
$4M^{(bb)}(\tau)=M_{c}^{(M)}+M_{0}(\tau)-\frac{M_{1}(\tau)+M_{1^{\prime}}(\tau)}{2}+M_{2}(\tau),$ (24a)
$2M^{(ab)}(\tau)=M_{c}^{(M)}-M_{0}(\tau)+\frac{M_{1}(\tau)-M_{1^{\prime}}(\tau)}{2}-M_{2}(\tau),$ (24b)
@@
for $j=c,0,1,1^{\prime},2$. It is easy to see that $M_{1}(\tau)=M_{1^{\prime}}(\tau)$, therefore we obtainEqs. (11) for the interferograms for both measurement schemes. Substituting the Gaussian photon pair spectral distribution (7) into Eq. (25) we get the final form of the interferogram terms (12).